Joe is ten years older than James. In eight years, twice Joe's age will equal three times James's age. How old is each at present?
Answer by JulietG(1812)

J = M + 10 (let M be James; why texts try to mess people up with similar names is beyond me).
2(J+8) = 3(M+8)
Distribute
2J + 16 = 3M + 24
Substitute the value of J from the first equation into the last.
2J + 16 = 3M + 24
2(M+10) + 16 = 3M + 24
Distribute
2M + 20 + 16 = 3M + 24
2M + 36 = 3M + 24
Subtract 2M from each side
36 = M + 24
Subtract 24 from each side
12 = M
.
If James is (12), then Joe is (22)
In 8 years, James will be 20 and Joe will be 30
Twice Joe's future age (30*2) is indeed 3x James' future age (20*3)